i have this problem:

m+6.5=9=-9m-130)
How do I Solve it?

Answer by checkley71(8403)

M+6.5=-9M-130
M+9M=-130-6.5
10M=-136.5
M=-136.5/10
M=-13.65 ANSWER.
PROOF:
-13.65+6.5=-9*-13.65-130
-7.15=122.85-130
-7.15=-7.15